Barrys BarBQue incurred the following costs: $1,400 for ribs, 45 hours of labor to cook the ribs at $10 per hour, $50 for seModel Magic Manufacturing reported the following year-end balances: Beginning work in process inventory, $35,000; beginning rSwifty Corporations accounting records reflect the following inventories: Dec. 31, 2020 Dec. 31, 2019 $410000 $360000 Raw maMarigold Corp. reported the following year-end information: $1080000 300000 Beginning work in process inventory Beginning rawSheridan Companys accounting records reflect the following inventories: Dec. 31, 2019 Dec. 31, 2020 $ 80000 $ 64000 Raw mate
Answer #1

1)Product costs are one which are used to manufacture or prepare a product . Includes direct material,direct labor, supplies and factory overhead

= 1400+(45*10)+50+150=2050

It is option D

2)Cost of goods manufactured= Cost of direct materials used+direct labor+total manufacturing overhead

=27600+20000+28900=$76,500

Option C

3) Work in process =Beginning work in process+cost of direct materials+ direct labor+ manufacturing overheads

=35000+(18000+510000-15000)+180000+75000=803,000

Option B

4) total manufacturing costs= cost of direct materials+ direct labor+ manufacturing overheads

cost of direct materials= Beg raw materials inv+purchases-ending raw material inv

=360000+900000-410000=$850,000

total manufacturing costs=850000+670000+640000=$2,160,000

Option A